OpenStudy (anonymous):

does x= 1/8 for (1/8)^(2x-5) = 4^(x+8)

OpenStudy (cwrw238):

plug x = 1/8 into the equation and see ; (1/8)^ (-19/4) = 19483 4 ^(8 1/8) = 77935 no

OpenStudy (anonymous):

No i need to find what x= though.

OpenStudy (anonymous):

How do you do it right

OpenStudy (cwrw238):

take logs of both sides (2x - 5) log 1/8 = (x+8) log 4 (2x-5) ------ = log 4 / log 1/8 = -2/3 (x+8) 2x - 5 = -(2/3)x - 16 /3 6x - `5 = -2x - `6 8x = -1 x = -1/8

OpenStudy (cwrw238):

you just had sign wrong this results checks out